# Task-Unaware Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

### v1

@@ 4.1 Data Retrieval
During deployment, we first retrieve the most relevant demonstrations from episodic memory $\mathcal{M}$ based on similarity to the current scenario. Due to the blurred task boundaries, some tasks share similar visual observations but differ in their task objectives, while others have similar goals but involve different backgrounds, objects, etc. To account for these variations, we compare both visual inputs from the workspace camera (Du et al., 2023) and task descriptions (de Masson D’Autume et al., 2019) using $L_{2}$ distances of their embeddings. The retrieval process follows a simple rule:
$\mathcal{D}_{R}=\alpha_{v}\cdot\mathcal{D}_{v}+\alpha_{l}\cdot\mathcal{D}_{l},$ (2)
where $\mathcal{D}_{R}$ is the weighted retrieval distance, $\mathcal{D}_{v}$ represents the distance between the embeddings of the scene observation from the workspace camera, and $\mathcal{D}_{l}$ depicts the distance between the task description embeddings. The parameters $\alpha_{v}$ and $\alpha_{l}$ control the relative importance of visual and language-based distances. Based on the distances $\mathcal{D}_{R}$, the most relevant demonstrations can be retrieved from $\mathcal{M}$.
@@ 4.2.1 Learn from Errors by Selective Weighting
To make the best use of the limited data, we enhance their utility by assigning weights to critical or vulnerable segments in each retrieved demonstration. Specifically, before testing, the robot performs several rollouts on the encountered task using the existing model trained during the lifelong learning phase. This procedure allows us to evaluate the model’s performance and identify any forgetting effects, akin to a preliminary quiz before the final exam (as illustrated in step 2, the reviewing phase in Figure 1).
When failed trajectories are identified, we compare each image in the retrieved demonstrations against all images from the failed trajectories using the $L_{2}$ distances of their embeddings. This comparison yields a distance vector for each demonstration, where each value represents the minimal distance between a demonstration frame and all images from the failed rollouts. This metric determines whether a particular frame has occurred during the rollout. Through this process, we identify the Separation Segment — frames in the demonstrations where the behavior deviates from what was executed during the failed rollouts (see Figure 3). Since these Separation Segments highlight behaviors that should have occurred but did not, we consider them vulnerable segments that contribute to the failure. We assign higher weights to these frames which will scale the losses during local adaptation. Detailed heuristics and implementation specifics are provided in Appendix A.4.
4.2.2 Adaptation with Fast Finetuning
Finally, we fine-tune the network’s parameters to better adapt to the current task using the retrieved demonstrations, focusing more on the difficult steps identified through selective weighting. Notably, the episodic memory $\mathcal{M}$ contains the same data used during training for experience replay and during deployment for local adaptation. No additional demonstrations are available to the robot at test time. Despite this limited data, our experiments demonstrate that the model can effectively recover learned skills and improve its performance across various tasks. Overall, the proposed weighted adaptation is formalized as follows:
@@ 4.2.2 Adaptation with Fast Finetuning
where $\tilde{N}$ is the number of retrieved demonstrations, $l_{n}$ is the length of demonstration $n$, and $w_{t,n}$ is the weight assigned to sample $t$ in demonstration $n$. The variables $o_{\leq t,n}$ and $a_{t,n}$ denote the sequence of observations up to time $t$ and the corresponding expert action, respectively, while $g_{n}$ is the goal description for demonstration $n$. The parameter $\mathbf{\theta}$ represents the network’s parameters before adaptation.
@@ 5 Experiments
We conduct a comprehensive set of experiments to evaluate the effectiveness of our proposed retrieval-based weighted local adaptation method for lifelong robot learning. Specifically, our experiments aim to address the following key questions:
@@
    Effect of Blurry Task Boundaries: How do blurry task boundaries influence the model’s performance and data retrieval during testing?
@@
    Advantages of Retrieval-Based Adaptation: Does retrieval-based weighted local adaptation enhance the robot’s performance across diverse tasks?
@@
    Impact of Selective Weighting: Is selective weighting based on rollout errors effective in improving task performance?
@@
    Generalizability: Can our method be applied to different memory-based lifelong robot learning approaches, serving as a paradigm that enhances the performance during test time by restoring previous knowledge and skills?
@@
    Robustness: Due to blurry task boundaries and retrieval imprecision, the retrieved demonstrations may not necessarily belong to the same task. How resilient is our method to inaccuracies in memory retrieval?
@@ 5.1.1 Benchmarks
We evaluate our proposed methods using LIBERO (Liu et al., 2024): libero_spatial, libero_object, libero_goal, and libero_different_scenes. These environments feature a variety of objects and layouts. The first three benchmarks all include $10$ distinct tasks, each with up to $50$ demonstrations collected in simulation with different initial states of objects and the robot. Specifically, libero_different_scenes is created from LIBERO’s provided LIBERO_90, which encompasses $20$ tasks from distinct scenes.
For each task, we paraphrased the assigned single goal description into diverse descriptions to obscure task boundaries. These enriched descriptions were generated by rephrasing the original task descriptions from the benchmark using a large language model provided by Phi-3-mini-4k-instruct Model (mini-4k instruct, 2024), ensuring consistent meanings while varying phraseology and syntax. Please see Section A.3 for more details.
@@ 5.1.2 Baselines
    Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017): A regularization-based approach that constrains updates to the network’s parameters to prevent catastrophic forgetting of previously learned tasks.
@@
    Experience Replay (ER) (Chaudhry et al., 2019): A core component of our training setup, ER utilizes stored episodic memory to replay past demonstrations, helping the model maintain previously acquired skills and mitigate forgetting. As a baseline, we evaluate the standalone performance of ER without additional retrieval-based adaptation techniques.
@@
    AGEM with Weighted Local Adaptation (AGEM-WLA): An extension of AGEM that incorporates weighted local adaptation during the testing phase, enhancing the model’s ability to adapt to specific tasks based on retrieved demonstrations. This allows us to assess the generalizability of our proposed method as a paradigm framework on other memory-based lifelong learning approaches.
@@
    PackNet (Mallya & Lazebnik, 2018): An architecture-based lifelong learning algorithm that iteratively prunes the network after training each task, preserving essential nodes while removing less critical connections to accommodate subsequent tasks. However, its pruning and post-training phases rely heavily on clearly defined task boundaries, making PackNet a reference baseline when task boundaries are well-defined.
@@ 5.1.3 Metrics
Our primary focus is on the success rate of task execution, as it is a crucial metric for manipulation tasks in interactive robotics. Consequently, we adopt the Average Success Rate (ASR) as our primary evaluation metric to address the challenge of catastrophic forgetting within the lifelong learning framework, evaluating success rates on three random seeds across all diverse tasks within the same benchmark.
5.1.4 Model, Training, and Adaptation
As illustrated in Figure. 2, our model utilizes pretrained encoders for visual and language inputs: R3M (Nair et al., 2022) for visual encoding, Sentence Similarity model (SS Model) (SentenceSimilarity, 2024) for language embeddings, and a trainable MLP-based network to encode proprioceptive inputs. Embeddings from ten consecutive time steps are processed through a transformer-based temporal encoder, with the resulting output passed to a GMM-based policy head for action sampling. Specifically, R3M, a ResNet-based model trained on egocentric videos using contrastive learning, captures temporal dynamics and semantic features from scenes, while Sentence Similarity Model captures semantic relationships in task descriptions, enabling the model to differentiate between various natural language instructions.
We trained the model sequentially on multiple tasks, with each task trained for $50$ epochs. For later tasks, we used experience replay, storing $8$ demonstrations per task from the episodic memory $\mathcal{M}$. Every 10 epochs, we evaluated the model on the current task over $20$ episodes, and the model with the highest ASR would be saved.
During testing, the agent performs 10 rollout episodes (referred to as a quiz phase) without adaptation, to assess initial performance and potential forgetting. We then retrieve the top $10\%$ most similar demonstrations from $\mathcal{M}$ using visual and language embeddings. Then, the model is fine-tuned for $20$ epochs with those weighted retrieved demonstrations. Finally, the adapted model is evaluated over $20$ episodes again (the final testing phase) to assess performance improvements. For experiments on all benchmarks, we train and test the models with three random seeds ($1$, $21$, and $42$) to reduce the impact of randomness.
@@ 5.2.1 Comparison with Baselines
To address Question 2, we compared our proposed method, Weighted Local Adaptation (ER-WLA), with all baseline approaches. As shown in Table 1, ER-WLA consistently outperforms baselines of EWC, AGEM, ER, and AGEM-WLA, which do not rely on clear task boundaries. By incorporating local adaptation during test time — our method mirrors how humans review and reinforce knowledge when it is partially forgotten — the continually learning robot could also regain its proficiency on previous tasks.
In contrast, PackNet serves as a reference method, as it requires well-defined task boundaries. However, as the number of tasks increases, the network’s trainable capacity under PackNet diminishes, leaving less flexibility for future tasks. This limitation becomes evident in the libero_different_scenes benchmark, which includes 20 tasks. PackNet’s success rate drops significantly for later tasks, resulting in poor overall performance and highlighting its constraints on plasticity compared with our proposed ER-WLA approach.
Additionally, when we applied WLA to the AGEM baseline (resulting in AGEM-WLA), it also improved its performance, demonstrating the effectiveness of our method as a paradigm for memory-based lifelong robot learning methods. These findings also support our conclusions regarding Question 4.
Table 1: Comparison with Baselines, the average success rates and standard deviations across various baselines are shown below. We provide PackNet’s performance on the right as a reference point for cases where task boundaries are accessible. Both EWC and vanilla AGEM demonstrate weak performance across all benchmarks, while ER performs better due to memory replay. Under our weighted local adaptation (WLA) paradigm, the WLA-enhanced versions of ER and AGEM show significant improvements over their vanilla counterparts, highlighting the effectiveness of WLA.
Method EWC AGEM AGEM-WLA ER ER-WLA PackNet
libero_spatial 0.0 $\pm$ 0.0 7.33 $\pm$ 14.25 35.83 $\pm$ 15.71 15.67 $\pm$ 13.50 39.83 $\pm$ 19.85 53.17 $\pm$ 24.72
libero_object 1.50 $\pm$ 3.26 27.17 $\pm$ 22.77 51.17 $\pm$ 24.13 56.50 $\pm$ 19.88 62.33 $\pm$ 18.69 73.77 $\pm$ 16.97
libero_goal 0.33 $\pm$ 1.83 10.83 $\pm$ 16.03 58.67 $\pm$ 25.93 52.33 $\pm$ 22.16 62.33 $\pm$ 28.75 66.33 $\pm$ 24.88
libero_different_scenes 2.58 $\pm$ 8.98 20.43 $\pm$ 25.55 41.75 $\pm$ 33.50 34.08 $\pm$ 28.55 45.17 $\pm$ 31.86 32.92 $\pm$ 44.03
@@ 5.2.2 Ablation Studies
In the first ablation, we evaluated the impact of selective weighting on libero_spatial, libero_object, and libero_goal benchmarks to demonstrate its importance for effective local adaptation. We compared two variants of our method: 1) ER-ULA, which applies uniform local adaptation without selective weighting, adapting retrieved demonstrations uniformly; 2) ER-WLA, which incorporates selective weighting during test-time adaptation. Both methods are trained with experience replay.
Since early stopping during local adaptation at test time is infeasible, and training can be unstable, particularly regarding manipulation success rates, we conducted adaptation using three different numbers of epochs — $15$, $20$, and $25$ — followed by final testing. The results, presented in Table 2, indicate that selective weighting enhances performance across different adaptation durations and various benchmarks, confirming our hypothesis in Question 3.
@@
(2) Due to this limitation, BERT is unable to retrieve the most relevant demonstrations (those most similar to the current task from the episodic memory $\mathcal{M}$). As a result, Retrieval-based WLA with BERT does not achieve optimal performance. These two findings address Question 1.
(3) Interestingly, from Figure 4 (c), despite BERT’s low Retrieval Accuracy (RA), if it attains a moderately acceptable rate (e.g., 0.375), the local adaptation using data retrieved based on BERT embeddings can still enhance model performance during test time. This demonstrates the robustness and fault tolerance of our proposed approach, further addressing Question 5.
Table 2: Ablation Study on Selective Weighting. This table presents the performance of success rates with uniform (ULA) and weighted (WLA) local adaptation across $15$, $20$, and $25$ epochs of adaptation under three random seeds, with evaluations conducted on all $10$ tasks within the benchmarks: libero_spatial, libero_object, and libero_goal. Compared to ULA, the weighted scheme significantly improves the method’s performance on most benchmarks.
Benchmark Method 15 Epochs 20 Epochs 25 Epochs Overall ASR (%)
ASR (%) ASR (%) ASR (%)
libero_spatial ER-ULA 35.33 $\pm$ 21.21 38.17 $\pm$ 14.76 38.16 $\pm$ 17.19 37.22 $\pm$ 17.77
ER-WLA 36.16 $\pm$ 18.55 39.83 $\pm$ 19.84 37.83 $\pm$ 17.70 37.94 $\pm$ 18.57
libero_object ER-ULA 57.83 $\pm$ 25.14 60.67 $\pm$ 22.96 58.00 $\pm$ 21.84 58.83 $\pm$ 23.13
ER-WLA 58.00 $\pm$ 22.35 62.33 $\pm$ 18.70 61.50 $\pm$ 24.36 60.61 $\pm$ 21.76
libero_goal ER-ULA 61.33 $\pm$ 28.43 62.00 $\pm$ 29.61 66.17 $\pm$ 27.22 63.17 $\pm$ 28.20
ER-WLA 62.83 $\pm$ 28.15 62.33 $\pm$ 28.76 67.50 $\pm$ 28.82 64.22 $\pm$ 28.35
((a)) PCA Visualization based on BERT
((b)) PCA Visualization based on SS Model
((c)) Bar Chart of Average Success Rates and Retrieval Accuracy across $10$ tasks
Figure 4: In Figure 4(a) and Figure 4(b), PCA is used to visualize the distribution of language embeddings of 3 tasks from BERT and SS, respectively. In Figure 4(c), SS model, which distinguishes task descriptions, has higher success rate and retrieval accuracy than BERT.

### v3

@@ 4.1 Data Retrieval
The proposed task-free lifelong learning algorithm retrieves relevant demonstrations from $\mathcal{M}$ based on similarity to the deployment scenario $S_{deploy}$. Besides, due to the blurry task boundaries, some tasks share similar visual observations but differ in their task objectives, while others have similar goals but involve different backgrounds, objects, etc. To account for these variations, the retrieval process compares both visual inputs from the workspace camera (Du et al., 2023) and language descriptions (de Masson D’Autume et al., 2019) using $L_{2}$ distances of their embeddings, following a simple rule:
$\mathcal{D}_{R}=\alpha_{v}\cdot\mathcal{D}_{v}+\alpha_{l}\cdot\mathcal{D}_{l},$
where $\mathcal{D}_{R}$ is the weighted retrieval distance, $\mathcal{D}_{v}$ represents the distance between the embeddings of the scene observation from the workspace camera, and $\mathcal{D}_{l}$ depicts the distance between the language description embeddings. The parameters $\alpha_{v}$ and $\alpha_{l}$ control the relative importance of visual and language-based distances. Based on the distances $\mathcal{D}_{R}$, the most relevant demonstrations can be retrieved from $\mathcal{M}$, as illustrated from Figure 3.
@@ 4.2.1 Learn from Errors by Selective Weighting
To make the best use of the limited data, we enhance their utility by assigning weights to critical or vulnerable segments in each retrieved demonstration. Specifically, before local adaptation, the robot performs several rollouts on the encountered scenario using the current model trained during the Lifelong Learning phase. This procedure allows us to evaluate the model’s performance and identify any forgetting effects (as illustrated in step 2, the Reviewing phase in Figure 1).
If a trial fails, we compare each image in the retrieved demonstrations against all images from the failed trajectories using $L_{2}$ distance of their embeddings. This comparison yields an Embedding Distance Matrix (EDM, shown in Figure 10) for each retrieved demonstration, where each value represents an embedding distance of a demonstration frame and an image from the failed rollout. This metric determines whether a particular frame has occurred during a failed rollout. Through this process, we identify the Separation Segment — frames in a demonstration where the failed rollout’s behavior starts to diverge from the demonstration (see Figure 4). Since these Separation Segments highlight expected behaviors that did not occur, we consider them critical points contributing to failure. We assign higher weights to these frames which will scale the losses during local adaptation. Detailed heuristics and implementation specifics are provided in Appendix A.4.
4.2.2 Local Adaptation with Fast Finetuning
Finally, we fine-tune the network’s parameters to better adapt to the deploying scenario $S_{deploy}$ using the retrieved demonstrations from $\mathcal{M}$, focusing more on the Separation Segments identified through selective weighting. Despite this limited data, our experiments demonstrate that the model can effectively recover learned knowledge and skills and improve robot’s performance across various tasks. Overall, the proposed weighted local adaptation is formalized as follows:
@@ 4.2.2 Local Adaptation with Fast Finetuning
where $\tilde{N}$ is the number of retrieved demonstrations, $l_{n}$ is the length of demonstration $n$, and $w_{t,n}$ is the weight assigned to sample $t$ in demonstration $n$. The variables $o_{\leq t,n}$ and $a_{t,n}$ denote the sequence of observations up to time $t$ and the corresponding expert action, respectively, while $g_{n}$ is the language description for demonstration $n$. The parameter $\mathbf{\theta}$ represents the network’s parameters before local adaptation.
@@ 5 Experiments
We conduct a comprehensive set of experiments to evaluate the effectiveness of RWLA for task-free lifelong robot learning. Specifically, our experiments aim to address the following key questions:
@@
    Effect of Blurry Task Boundaries: How do the blurry task boundaries influence the model’s performance and data retrieval during testing?
@@
    Advantages of RWLA: Does the proposed approach enhance the robot’s performance across diverse tasks?
@@
    Impact of Selective Weighting: Is selective weighting based on rollout failures effective?
@@
    Generalizability: Can our method be applied to different memory-based lifelong robot learning approaches, serving as a paradigm that enhances performance?
@@
    Robustness: How robust is our approach to imperfect demonstration retrieval, particularly when ambiguous task boundaries cause retrieved examples to mismatch the deploying scenario?
@@ 5.1.1 Benchmarks
We evaluate our proposed methods using LIBERO benchmarks (Liu et al., 2024): libero_spatial, libero_object, libero_goal, and libero_different_scenes. These environments feature a variety of task goals, objects, and layouts. The first three benchmarks all include $10$ distinct task goals (e.g., “Put the bottle into the basket.”, “Open the middle drawer of the cabinet.”), each with up to $50$ demonstrations collected from sampled simulation scenarios with different initial states of objects and the robot. Specifically, libero_different_scenes is created from LIBERO’s LIBERO_90, which encompasses $20$ tasks from distinct scenes.
We paraphrased the assigned single task goal into diverse language descriptions to obscure task boundaries (See Figure 7). These enriched language descriptions were generated by rephrasing the original task goal from the benchmark using a large language model provided by Phi-3-mini-4k-instruct Model (mini-4k instruct, 2024), ensuring consistent meanings while varying phraseology and syntax. Please see Section A.3 for more details.
@@ 5.1.2 Baselines
    Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017): A regularization-based approach that relies on task boundaries and restricts network parameters’ updates to prevent catastrophic forgetting of previously learned tasks.
@@
    Experience Replay (ER) (Chaudhry et al., 2019): A core component of our training setup, ER utilizes a storage memory to replay past demonstrations, helping the model maintain previously acquired skills and mitigate forgetting.
@@
    AGEM-RWLA: An extension of AGEM that incorporates RWLA before policy deployment, enhancing the model’s ability to adapt to specific scenarios. This allows us to assess the generalizability of our proposed method as a paradigm framework on other memory-based lifelong learning approaches.
@@
    PackNet (Mallya & Lazebnik, 2018): An architecture-based lifelong learning algorithm that iteratively prunes the network after training each task, preserving essential nodes while removing less critical connections to accommodate subsequent tasks. However, its pruning and post-training phases rely heavily on clearly defined task IDs, making PackNet a reference baseline when the IDs are well-defined.
Table 1: Comparison with Baselines. The Average Success Rates (ASR, %) and standard deviations (STD, %) across various baselines are shown below. We provide PackNet’s performance as a reference point for cases where task IDs are accessible. Both EWC and vanilla AGEM demonstrate weak performance across all benchmarks. Under our Retrieval-based Weighted Local Adaptation (RWLA) paradigm, both ER and AGEM show significant improvements over their vanilla counterparts, highlighting the effectiveness of RWLA.
[…]
EWC PackNet AGEM AGEM-RWLA ER ER-RWLA (ours)
libero_spatial 0.0 $\pm$ 0.0 53.17 $\pm$ 7.25 7.33 $\pm$ 1.61 35.83 $\pm$ 9.39 15.67 $\pm$ 6.66 39.83 $\pm$ 8.13
libero_object 1.50 $\pm$ 0.50 73.67 $\pm$ 1.89 27.17 $\pm$ 4.65 51.17 $\pm$ 3.62 56.50 $\pm$ 3.97 62.33 $\pm$ 1.76
libero_goal 0.33 $\pm$ 0.58 66.33 $\pm$ 4.19 10.83 $\pm$ 3.55 58.67 $\pm$ 3.21 52.33 $\pm$ 5.20 62.33 $\pm$ 4.93
libero_different_scenes 2.58 $\pm$ 1.14 32.92 $\pm$ 0.63 20.43 $\pm$ 1.17 41.75 $\pm$ 5.66 34.08 $\pm$ 2.47 45.17 $\pm$ 0.38
Table 2: Ablation Study on Selective Weighting. This table presents ASR (%) and their STD (%) for uniform (RULA) and weighted (RWLA) local adaptation across $15$, $20$, and $25$ epochs of adaptation under three random seeds, with evaluations conducted on all $10$ tasks within the benchmarks: libero_spatial, libero_object, and libero_goal. Compared to RULA, selective weighting scheme improves the method’s performance on most benchmarks.
Benchmark Method 15 Epochs 20 Epochs 25 Epochs Overall ASR
libero_spatial RULA 35.33 $\pm$ 9.07 38.17 $\pm$ 5.06 38.17 $\pm$ 3.33 37.22 $\pm$ 5.64
RWLA 36.17 $\pm$ 8.33 39.83 $\pm$ 8.13 37.83 $\pm$ 3.75 37.94 $\pm$ 6.32
libero_object RULA 57.83 $\pm$ 9.07 60.67 $\pm$ 3.01 58.00 $\pm$ 4.92 58.83 $\pm$ 5.55
RWLA 58.00 $\pm$ 4.44 62.33 $\pm$ 1.76 61.50 $\pm$ 1.32 60.61 $\pm$ 3.18
libero_goal RULA 61.33 $\pm$ 3.69 62.00 $\pm$ 2.60 66.17 $\pm$ 5.58 63.17 $\pm$ 4.24
RWLA 62.83 $\pm$ 7.65 62.33 $\pm$ 4.93 67.50 $\pm$ 1.80 64.22 $\pm$ 5.26
@@ 5.1.3 Metrics
We focus on the success rate of task execution, as it is a crucial metric for manipulation tasks in interactive robotics. Consequently, we adopt the Average Success Rate (ASR) as our primary evaluation metric to address the challenge of catastrophic forgetting within the lifelong learning framework, evaluating success rates on three random seeds across all diverse tasks within each benchmark.
5.1.4 Model, Training, and Evaluation
As illustrated in Figure 2, our model utilizes pretrained encoders for visual and language inputs: R3M (Nair et al., 2022) for visual encoding, Sentence Similarity model (SS Model) (SentenceSimilarity, 2024) for language embeddings, and a trainable MLP-based network to encode proprioceptive inputs. Embeddings from ten consecutive time steps are processed through a transformer-based temporal encoder, with the resulting output passed to a GMM-based policy head for action sampling. Specifically, R3M, a ResNet-based model trained on egocentric videos using contrastive learning, captures temporal dynamics and semantic features from scenes, while Sentence Similarity Model captures semantic meanings in language descriptions, enabling the model to differentiate between various instructions.
To standardize the comparisons with baseline lifelong robot learning algorithms in LIBERO benchmarks, the model first undergoes a Lifelong learning phase, where it is trained sequentially on demonstrations from $10$ or $20$ tasks, depending on the specific benchmark, with each task trained for $50$ epochs. A small number of demonstrations from each task is stored in $\mathcal{M}$, playing a dual-use for experience replay and RWLA. Every $10$ epochs, we check the model’s performance and save the version that achieves the highest Success Rate to prevent over-fitting.
After training on all tasks sequentially, we conduct reviewing and testing on various scenarios sampled from each task for comprehensive analysis. During the reviewing stage, we firstly evaluate potential forgetting by having the agent perform $10$ rollout episodes on the deployment scenario $\mathcal{S}_{deploy}$. We then retrieve the most similar demonstrations from $\mathcal{M}$ and fine-tune the model for only $20$ epochs using the retrieved demonstrations with selective weighting. Finally, we deploy the adapted model for $20$ episodes—the testing phase—to assess performance improvements. All training, local adaptation, and testing in the benchmarks are conducted using three random seeds ($1$, $21$, and $42$) to reduce the impact of randomness.
@@ 5.2.1 Comparison with Baselines
To address Question 2, we compared RWLA, with all baseline approaches. As shown in Table 1, ER-RWLA consistently outperforms baselines of EWC, AGEM, ER, and AGEM-RWLA. By incorporating local adaptation before policy deployment — our method mirrors how humans review and reinforce knowledge when it is partially forgotten — the continually learning robot could also regain its proficiency on previous tasks.
In contrast, PackNet serves as a reference method, as it requires explicit task IDs. Noteworthy, as the number of tasks increases, the network’s trainable capacity under PackNet diminishes, leaving less flexibility for future tasks. This limitation becomes evident in the libero_different_scenes benchmark, which includes 20 tasks, see Table 8. PackNet’s success rate drops significantly for later tasks, resulting in poor overall performance and highlighting its constraints on plasticity compared with the proposed ER-RWLA.
Additionally, when we applied RWLA to the AGEM baseline (AGEM-RWLA), it also improved its performance, demonstrating the effectiveness of our method as a paradigm for memory-based lifelong robot learning methods. These findings support our conclusions regarding Question 4.
(a) PCA based on BERT
(b) PCA based on SS Model
(c) Bar Chart of Average Success Rates and Retrieval Accuracy across $10$ tasks
Figure 5: In Figures 5(a) and 5(b), PCA is used to visualize the distribution of language embeddings of 3 tasks from BERT and SS, respectively. In Figure 5(c), the SS model, which distinguishes task descriptions, has higher average success rates (ASR) and retrieval accuracy (RA) than BERT. The error bars represent the standard deviations of ASR and RA for each task over 20 repetitions with 3 random seeds.
@@ 5.2.2 Ablation Studies
In the first ablation, we evaluated the impact of selective weighting on libero_spatial, libero_object, and libero_goal benchmarks to demonstrate its importance for effective local adaptation. We compared a variant of RWLA: RULA, which applies uniform local adaptation without selective weighting, adapting retrieved demonstrations uniformly. Both methods are trained with ER.
Since early stopping during local adaptation at test time is infeasible, and training can be unstable, particularly regarding manipulation success rates, we conducted RWLA using three different numbers of epochs — $15$, $20$, and $25$. The results presented in Table 2, indicate that selective weighting enhances performance across different adaptation durations and various benchmarks, addressing Question 3.
@@
(2) Due to this limitation, BERT is unable to retrieve the most relevant demonstrations (those most similar to the current scenario from the storage memory $\mathcal{M}$). As a result, RWLA with BERT does not achieve optimal performance. These two findings address Question 1.
(3) Interestingly, from Figure 5 (c), despite BERT’s low Retrieval Accuracy (RA), if it attains a moderately acceptable rate (e.g., 0.375), the RWLA based on BERT embeddings can still enhance model performance. This demonstrates the robustness and fault tolerance of our proposed approach, further addressing Question 4 and 5.